OpenStudy (anonymous):

Which Expression is equivalent to 2b - 3a +b +a? A. b-2a B. b-4a C. 3b -4a D. 3b-2a I think its d but I don't get where the extra +a and +b would go if that was the case... Can anyone explain?

OpenStudy (anonymous):

2b - 3a + b + a combine like terms: 2b + b = 3b -3a + a = -2a leaving us with : 3b - 2a your right...it is D

OpenStudy (anonymous):

when there is just a letter - such as "a", it means "1a"

OpenStudy (anonymous):

Whenever there is a letter by itself, imagine there is a one in front of it.

OpenStudy (anonymous):

Yeah I know that but I was thinking it would add the 2a^2 but thanks :D

OpenStudy (anonymous):

glad to help....when adding like terms, the variables stay the same.. 3a + 2a = 5a 15a - 12a = 3a well...you see what I mean